# OW-Rep: Open World Object Detection with Instance Representation Learning

### v2

@@ 4 Experiments
For more comprehensive understanding, please refer to the supplementary material (Section  A and  B).
@@ 4.1 Datasets
We conducted experiments using splits of the MS-COCO [28] and Pascal VOC [9] datasets. In the OWOD split [18], all 80 classes are divided into four sequential tasks, $\{T_{1},T_{2},T_{3},T_{4}\}$, to simulate an incremental learning setting where each task contains 20 classes. During training for task $T_{c}$, only the 20 newly introduced classes in $T_{c}$ are labeled. Trained on this dataset, the detector is required to detect both the known classes from previously encountered tasks $\{T_{t}\mid t\leq c\}$ and the unknown classes from future tasks $\{T_{t}\mid t>c\}$. Additionally, a few exemplars from previously known classes are stored in a replay buffer to mitigate catastrophic forgetting, which are used to fine-tune the detector at the end of each task [50]. By repeating this process, the detector continuously discovers and learns new object classes, gradually adapting to the open-world.
The OWOD split allows unknown classes to appear in training images as unlabeled objects. However, it does not account for real-world scenarios where entirely novel objects that were never seen during training may appear during deployment. To better reflect these realistic conditions and to evaluate the detector’s ability to detect and represent such unknown-unknown objects, we propose a new Unknown-Unknown split. This split builds upon the OWDETR split [13], which strictly separates superclasses across tasks. In addition, we exclude all images containing unknown classes from the training set, ensuring the detector has no prior exposure to these objects. This exclusion enables a more rigorous evaluation of the detector’s ability to learn a generalizable feature embedding space for completely novel objects.
Table 2: Quantitative results on Unknown-Unknown split. We compared our method with the other OWOD methods that learn instance embeddings, reporting performance on Unknown Recall, Recall@1, and known mAP. While existing methods show decreased performance as unknown-unknowns are never encountered during training, our method shows large improvement in both unknown object detection and feature embedding quality. This demonstrates that our method learns robust and generalizable instance embedding space.
@@
PROB[50] 21.04 3.32 67.65 29.85 7.81 44.80 33.37 12.62 39.65 38.70
OSODD[47] 21.04 3.78 67.65 29.85 8.53 44.80 33.37 11.56 39.65 38.70
UC-OWOD[46] 14.80 2.41 28.52 19.47 4.87 25.04 21.98 7.43 26.96 38.07
RNCDL[11] 21.91 1.72 66.73 27.81 2.40 42.51 32.17 4.96 39.17 37.08
@@ 4.2 Experimental Setup
Implementation Details. We utilized PROB [50] for the open-world object detector, which employs the deformable DETR [48] with a DINO-pretrained [3] Resnet-50 FPN backbone [13]. The model uses $100$ input queries, each with a dimension of ${d=256}$. Training is performed for 41 epochs, with the losses $\mathcal{L}_{b,\text{unk}}$ and $\mathcal{L}_{et}$ applied after the initial warmup period, using weight coefficients of ${\alpha=0.1}$, ${\beta=1.0}$, respectively. The batch size is set to 32, and the AdamW optimizer is employed with a learning rate of $3e^{-4}$.
For unknown box refinement, we employed the ViT-H SAM model [22] to generate masks, using the top ${k=10}$ proposals with the highest objectness scores as prompts. For embedding transfer, we used the ViT-L DINOv2 model [34] with a feature dimension of ${D=1024}$. The remaining hyperparameters are set as follows: ${\kappa=0.5}$, ${\sigma=1.0}$, ${\delta=1.0}$.
Evaluation Metrics. For known classes, the mean Average Precision (mAP) is used for evaluation. However, for unknown objects, false positives become unreliable, as not every object in the dataset is labeled. Therefore, following previous works [30, 13, 50], the Unknown-Recall (U-Recall) is used as a metric for unknown object detection.
To evaluate the quality of object embeddings, we first assign each ground truth object to the instance embedding of the predicted proposal with the highest overlap. Using these ground truth labels and corresponding features, we compute Recall@$K$ as a performance metric [21], which measures how many objects have at least one sample from the same class among their $K$-nearest neighbors in the learned embedding space. When computing Recall@$K$, we assume that undetected ground truth objects have no valid corresponding samples in their $K$-nearest neighbors, as the detector failed to generate embeddings for them.
Comparison Methods. We compared our method with several other approaches to demonstrate its effectiveness in detecting and extracting instance embeddings for unknown objects. For the base model, PROB [50], the detector’s query embeddings are directly used as instance embeddings. We also evaluated methods that enhance feature quality through self-supervision. OSODD [47] augments cropped proposals from the pretrained and frozen detector and trains a separate proposal embedding network using contrastive learning [15]. In contrast, UC-OWOD [46] and RNCDL [11] directly optimize the detector’s instance embeddings. UC-OWOD applies contrastive learning between instance embeddings of high-confidence unknown proposals, where only highly similar or highly dissimilar pairs are used as positive or negative samples. Meanwhile, RNCDL clusters all instance embeddings based on similarity and treats all embeddings within the same cluster as positive pairs. All methods are reimplemented based on the PROB framework.
@@ 4.3 Experimental Results
Table 4: Ablation results for base Open World Object Detector. Applying both the proposed modules (Ours) to other open-world object detectors, including OW-DETR [13] and Hyp-OW [8], our module improves Unknown Recall and Recall@1. These results demonstrate that the proposed modules consistently enhance instance embedding quality and unknown object detection across different backbones.
@@
To evaluate the impact of our method on instance embedding quality, we visualize the instance embedding space using t-SNE, as shown in Fig. 5. Unlike PROB, where embeddings are mixed and lack clear separation, our method produces a well-structured feature space with distinct class-wise clusters. Additionally, semantically similar animal classes appear closer together, while unrelated categories remain separate. This shows that our method effectively captures semantic relationships within the feature space.
@@
Table 5: Quantitative results for open-world tracking. Ours outperforms others in both unknown detection performance and the association of detections between frames. A-Accuracy represents the Association Accuracy [30].
@@
PROB [50] 48.67 8.49 19.72
RNCDL [11] 35.45 8.30 16.74
@@ 4.4 Open-World Tracking
To validate the applicability of our method, which detects both known and unknown objects while extracting semantically rich features, we apply it to the downstream task of open-world tracking [30]. In tracking scenarios, objects often become occluded or undergo deformation, making robust features essential for accurate re-identification. By detecting and extracting semantically rich embeddings of both known and unknown objects, our method helps the tracker associate objects more reliably under open-world conditions. We adopt DeepSORT [44] as the tracker, which matches proposals across adjacent frames by combining the IoU of bounding boxes with the similarity between the detector’s instance embeddings. The experiment uses a subset of the TAO-OW dataset [30] containing objects with dynamic motion. Evaluation focuses only on unknown objects, following the OWTA metric [30], which combines unknown detection recall with association accuracy.
@@ 5 Conclusions
This paper extends open-world object detection by simultaneously detecting objects and extracting semantically rich features under open-world conditions. By distilling both predictions and features of Vision Foundation Models, the proposed method accurately localizes unknown objects and captures the nuanced relationships through instance embeddings. Extensive experiments show its effectiveness in learning a semantically rich and generalizable feature space for instance embeddings, expanding the applicability of open-world object detectors. Future work will explore online learning techniques to further refine the feature space during inference.
Appendix A Additional Experiments Material
A.1 Proposed Data Split
The Unknown-Unknown split is built upon the OWDETR split, which was designed to minimize the presence of visually similar objects within the same data split. While the OWDETR split includes unknown class objects in the training set without labeling them, our proposed split goes a step further by entirely removing images containing unknown class objects from the training data. The details presented in Table  S1. Consequently, the model has no prior exposure to these unknown-unknown objects. This data split enables a more rigorous evaluation of whether an open-world object detector can effectively identify truly unknown objects and capture their inter-proposal relationships by learning semantically rich feature representations.
@@ A.1 Proposed Data Split
A.2 Motivation for Evaluation Metrics
@@
A.3 Additional Implementation Details
During the initial training phase, only $L_{o}$, $L_{c}$, and $L_{b}$ are optimized. The Embedding Transfer Module and Unknown Box Refine Module are introduced after 14 and 35 epochs, respectively. For the Unknown Box Refine Module, four regular grid points are sampled from each input unknown proposal, and used as prompts to generate a instance mask. The refined unknowns are defined as the smallest bounding box containing this mask. Bounding boxes are used for training only if their IoU before and after refinement is 0.5 or higher, minimizing performance degradation on known classes.
Appendix B Additional Results
@@
PROB 18.84 4.42 58.98
PROB + L1 Distill 31.19 8.69 59.20
Ours 31.54 11.40 59.12

### v3

@@ 4 Experiments
For more comprehensive understanding, please refer to the supplementary material (Sections A, B, C, D , and E).
@@ 4.1 Datasets
We conducted experiments using splits of the MS-COCO [lin2014microsoft] and Pascal VOC [everingham2010pascal] datasets. In the OWOD split [joseph2021towards], all 80 classes are divided into four sequential tasks, $\{T_{1},T_{2},T_{3},T_{4}\}$, to simulate an incremental learning setting where each task contains 20 classes. During training for task $T_{c}$, only the 20 newly introduced classes in $T_{c}$ are labeled. Trained on this dataset, the detector is required to detect both the known classes from previously encountered tasks $\{T_{t}\mid t\leq c\}$ and the unknown classes from future tasks $\{T_{t}\mid t>c\}$. Additionally, a few exemplars from previously known classes are stored in a replay buffer to mitigate catastrophic forgetting, which are used to fine-tune the detector at the end of each task [zohar2023prob]. By repeating this process, the detector continuously discovers and learns new object classes, gradually adapting to the open world.
The OWOD split allows unknown classes to appear in training images as unlabeled objects. However, it does not account for real-world scenarios where entirely novel objects that were never seen during training may appear during deployment. To better reflect these realistic conditions and to evaluate the detector’s ability to detect and represent such unknown-unknown objects, we propose a new Unknown-Unknown split. This split builds upon the OWDETR split [gupta2022ow], which strictly separates superclasses across tasks. In addition, we exclude all images containing unknown classes from the training set, ensuring the detector has no prior exposure to these objects. This exclusion enables a more rigorous evaluation of the detector’s ability to learn a generalizable feature embedding space for completely novel objects.
Table 2: Quantitative results on Unknown-Unknown split. We compared our method with the other approaches that simultaneously detect unknown objects and learn instance embeddings. While existing methods show decreased performance as unknown-unknowns are never encountered during training, our method shows large improvement in both unknown object detection and feature embedding quality. This demonstrates that our method learns robust and generalizable instance embedding space.
@@
PROB[zohar2023prob] 21.04 3.32 67.65 29.85 7.81 44.80 33.37 12.62 39.65 38.70
OSODD[zheng2022towards] 21.04 3.78 67.65 29.85 8.53 44.80 33.37 11.56 39.65 38.70
UC-OWOD[wu2022uc] 14.80 2.41 28.52 19.47 4.87 25.04 21.98 7.43 26.96 38.07
RNCDL[fomenko2022learning] 21.91 1.72 66.73 27.81 2.40 42.51 32.17 4.96 39.17 37.08
@@ 4.2 Experimental Setup
Implementation Details. We utilized PROB [zohar2023prob] for the open-world object detector, which employs deformable DETR [zhu2020deformable] with a DINO-pretrained [caron2021emerging] Resnet-50 FPN backbone [gupta2022ow]. The model uses $100$ input queries, each with an embedding dimension of d=256{d=256}. Training runs for 41 epochs, with the losses $\mathcal{L}_{b,\text{unk}}$ and $\mathcal{L}_{et}$ applied after the initial warmup period, using weight coefficients of $\alpha=0.1$}, $\beta=1.0$}. The batch size is set to 32, and the AdamW optimizer is employed with a learning rate of $3e^{-4}$.
For unknown box refinement, we employed the ViT-H SAM model [kirillov2023segment] to generate masks, using the top k=10{k=10} proposals with the highest objectness scores as prompts. For embedding transfer, we used the ViT-L DINOv2 model [oquab2023dinov2] with a feature dimension of D=1024{D=1024}. The remaining hyperparameters are set as follows: $\kappa=0.5$}, $\sigma=1.0$}, $\delta=1.0$}.
Evaluation Metrics. For known classes, the mean Average Precision (mAP) is used for evaluation. However, for unknown objects, false positives become unreliable, as not every object in the dataset is labeled. Therefore, following previous works [liu2022opening, gupta2022ow, zohar2023prob], the Unknown-Recall (U-Recall) is used as a metric for unknown object detection.
To evaluate the quality of object embeddings, we first assign each ground truth object to the instance embedding of the predicted proposal with the highest overlap. Using these ground truth labels and corresponding features, we compute Recall@$K$ as a performance metric [kim2021embedding], which measures how many objects have at least one sample from the same class among their $K$-nearest neighbors in the learned embedding space. When computing Recall@$K$, we assume that undetected ground truth objects have no valid corresponding samples in their $K$-nearest neighbors, as the detector failed to generate embeddings for them.
Comparison Methods. We compared our method with several other approaches to demonstrate its effectiveness in detecting and extracting instance embeddings for unknown objects. For the base model, PROB [zohar2023prob], the detector’s query embeddings are directly used as instance embeddings. We also evaluated methods that enhance feature quality through self-supervision. OSODD [zheng2022towards] augments cropped proposals from the pretrained frozen detector and trains a separate proposal embedding network using contrastive learning [he2020momentum]. In contrast, UC-OWOD [wu2022uc] and RNCDL [fomenko2022learning] directly optimize the detector’s instance embeddings. UC-OWOD applies contrastive learning between instance embeddings of high-confidence unknown proposals, using only highly similar or highly dissimilar pairs as positive or negative samples. Meanwhile, RNCDL clusters all instance embeddings based on similarity and treats all embeddings within the same cluster as positive pairs. All methods are reimplemented based on the PROB framework.
@@ 4.3 Experimental Results
We first compare our method with existing approaches that jointly perform unknown object detection and instance feature learning. We then present ablation studies on each module and baseline detector to demonstrate the effectiveness and compatibility of our approach.
@@
Table 4: Ablation results for baseline Open-World Object Detector. Applying our modules to other open-world object detectors, including OW-DETR [gupta2022ow] and Hyp-OW [doan_2024_HypOW], our module improves Unknown Recall and Recall@1. This demonstrates that our method consistently enhances instance embedding quality and unknown object detection across different backbones. (see Sec B.1. in the supplement)
@@
To evaluate the impact of our method on instance embedding quality, we visualize the instance embedding space using t-SNE, as shown in Fig. S3. Unlike PROB, where embeddings are mixed and lack clear separation, our method produces a well-structured feature space with distinct class-wise clusters. Additionally, semantically similar animal classes appear closer together, while unrelated categories remain separate. This shows that our method effectively captures semantic relationships within the feature space.
@@
Table 5: Quantitative results for open-world tracking. Ours outperforms others in both unknown detection performance and the association of detections between frames. A-Accuracy represents the Association Accuracy [liu2022opening].
@@
PROB [zohar2023prob] 48.67 8.49 19.72
RNCDL [fomenko2022learning] 35.45 8.30 16.74
@@ 4.4 Open-World Tracking
To validate the applicability of our method, which detects both known and unknown objects while extracting semantically rich features, we apply it to the downstream task of open-world tracking [liu2022opening]. In tracking scenarios, objects often become occluded or undergo deformation, making robust features essential for accurate re-identification. By detecting and extracting semantically rich embeddings of both known and unknown objects, our method helps the tracker associate objects more reliably under open-world conditions. We adopt DeepSORT [wojke2017simple] as the tracker, which matches proposals across adjacent frames by combining the IoU of bounding boxes with the similarity between the detector’s instance embeddings. The experiment uses a subset of the TAO-OW dataset [liu2022opening] containing objects with dynamic motion. Evaluation focuses only on unknown objects, following the OWTA metric [liu2022opening], which combines unknown detection recall with association accuracy.
@@ 5 Conclusions
This paper presents a method to extend open-world object detection by simultaneously detecting objects and extracting semantically rich features under open-world conditions. By distilling both predictions and features of Vision Foundation Models, the proposed method accurately localizes unknown objects and captures nuanced relationships between instances through instance embeddings. Extensive experiments demonstrate the effectiveness of our approach in learning a semantically rich and generalizable feature space for instance embeddings, expanding the applicability of open-world object detectors. To further enhance the real-world applicability, future work will explore online learning techniques to refine the feature space during inference.
6 Acknowledgments
This work was partly supported by Institute for Information & communications Technology Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST) and the Agency For Defense Development Grant funded by the Korean Government.
Supplementary Materials
Contents
Appendix A Problem Setting: OWOD vs OVOD
Figure S1: Limitations of OVOD with text prompts. Grounding DINO 1.6 [ren2024grounding] is utilized for the experiment. (a) When the prompt is given as Cat, Monitor, Object, the model detects the cat and the monitor, but fails to recognize the keyboard. (b) When the prompt includes Keyboard, the model successfully detects it. This demonstrates that OVOD methods are heavily dependent on explicit text prompts, and objects not included in the prompt cannot be detected.
@@ Appendix A Problem Setting: OWOD vs OVOD
To clarify the scope of our work, we distinguish between the problem settings of Open Vocabulary Object Detection (OVOD) and Open-World Object Detection (OWOD), and highlight their complementary nature. OVOD focuses on detecting objects in an image based on given text prompts [gu2021open, kuo2022f, wu2023aligning, li2023distilling, wang2023open, wang2023learning, zang2022open, li2022grounded, liu2024grounding, ren2024grounding]. These methods typically leverage pretrained language models to fuse vision and language modalities, grounding the detector through explicit textual supervision. In contrast, OWOD, including our proposed approach, operates without any external input beyond the image itself. The goal is to detect both known objects (seen during training) and unknown objects (unseen categories), which can then be relabeled by human annotators and incrementally incorporated into the model to expand its knowledge.
While OVOD naturally learns semantics of each object during grounding, previous OWOD methods focus on detecting unknown objects, without explicitly learning their semantics. Some OVOD models like Grounding DINO [liu2024grounding] and GLIP [li2022grounded] are trained on vast, diverse detection datasets far exceeding COCO, with the primary objective of enabling zero-shot detection through text prompts. Consequently, their framework does not inherently address scenarios that lack text prompts or require incremental learning, which are central to the OWOD paradigm.
Our setting assumes that the set of objects to be detected is not known in advance and that corresponding text prompts are unavailable. This reflects a realistic open-world scenario where detectors inevitably encounter unexpected objects. In such cases, requiring text prompts for every possible object is impractical. Consequently, the fundamental difference between OWOD and OVOD lies in whether object detection depends on pre-specified text prompts.
Due to this difference in formulation, we do not directly compare our method with OVOD approaches. OVOD is suitable for cases where explicit prompts are provided, whereas OWOD addresses scenarios where such prompts cannot be assumed. For example, as shown in Fig S1, Grounding DINO can detect a keyboard only when given the corresponding prompt. In deployment contexts such as autonomous driving, detectors must recognize previously unseen objects without relying on textual input—conditions that naturally fall under the OWOD setting rather than OVOD.
Moreover, a central objective of OWOD is not only to detect unknown objects but also to distinguish them from known classes and enable incremental learning, where newly identified categories can be incorporated into the model over time. These differences emphasize that OWOD and OVOD are complementary but fundamentally distinct research directions, rather than directly comparable approaches.
Appendix B Additional Experiments Material
B.1 Proposed Data Split
The Unknown-Unknown split is built upon the OWDETR split, which was designed to minimize the presence of visually similar objects within the same data split. While the OWDETR split includes unknown class objects in the training set without labeling them, our proposed split goes a step further by entirely removing images containing unknown class objects from the training data. The details presented in Table  S1. To illustrate, consider the Task 1 setting of the OWDETR split, where the model is trained on known classes such as animals, persons, and vehicles. In this setup, the training set often includes images where known objects co-occur with instances of unknown, future-task classes—for example, a person in a kitchen with a donut or a sink. The standard protocol discards only the annotations for donut and sink but still uses the image for training. In contrast, our Unknown-Unknown split excludes the entire image from the training data, thereby preventing the model from being visually exposed to any unknown objects. Consequently, the model has no prior exposure to these unknown-unknown objects. This data split enables a more rigorous evaluation of whether an open-world object detector can effectively identify truly unknown objects and capture their inter-proposal relationships by learning semantically rich feature representations.
Figure S2: A failure case of the SAM-based box refinement. Left: The initial bounding boxes predicted by the detector. Right: The refined boxes generated from SAM’s masks. Boxes of the same color correspond to the proposals before and after refinement. The green box highlights a significant discrepancy that can occur without IoU thresholding, underscoring its importance.
B.2 Motivation for Evaluation Metrics
@@
B.3 Additional Implementation Details
During the initial training phase, only $L_{o}$, $L_{c}$, and $L_{b}$ are optimized. The Embedding Transfer Module and Unknown Box Refine Module are introduced after 14 and 35 epochs, respectively. For the Unknown Box Refine Module, four regular grid points are sampled from each input unknown proposal, and used as prompts to generate a instance mask. The refined unknowns are defined as the smallest bounding box containing this mask. Bounding boxes are used for training only if their IoU before and after refinement is 0.5 or higher, minimizing performance degradation on known classes. This ensures the detector from training with wrong supervisions, such as depicted in Fig S2
Appendix C Ablation and Component Analysis
C.1 Additional Ablation for Baseline Detector.
Our approach is designed as modular components that can be integrated into any baseline open-world object detector. We primarily adopt PROB as the baseline detector to highlight the effectiveness of our method in improving both unknown object detection and instance-level feature representation. Additionally, we validate its general applicability by conducting experiments with multiple OWOD baselines Table 4 and Table S2.
Table S2: Ablation results for base Open World Object Detector. U-Recall denotes the unknown detection recall. Incorporating our module enhances both instance embedding quality and unknown detection performance.
[…]
CAT [ma2023cat] 9.06 3.07 59.33
CAT+Ours 11.67 4.81 59.75
OrthoDet [sun2024exploring] 21.56 11.84 60.83
OrthoDet+Ours 29.74 13.35 59.41
C.2 Comparison with using SAM and DINOv2 features for ETM
Table S3 presents results of distilling feature similarity from SAM and DINOv2. Although embedding transfer can use any source model, we use DINOv2 for its superior feature learning performance.
Table S3: Results using SAM and DINOv2 features for embedding transfer. U-Recall denotes the unknown detection recall. Distilling DINOv2 features further improves feature quality.
[…]
Ours (w. SAM) 30.63 10.38 59.38
Ours (w. DINOv2) 30.56 11.69 58.89
C.3 Ablation Studies of Distillation Method
[…]
PROB 18.84 4.42 58.98
PROB + L1 Distill 31.19 8.69 59.20
Ours 31.54 11.40 59.12
@@ C.3 Ablation Studies of Distillation Method
C.4 Ablation Studies of Module under Unknown-Unknown split
C.5 Ablation Studies of each Module’s Hyperparameters
We performed a comprehensive hyperparameter ablation study to validate the robustness of the ETM and URM modules within the OWOD split. The experimental results are detailed in Table S6. Our model demonstrates insensitivity to variations in each hyperparameter. Furthermore, we observed that varying parameters like $k$ and $\kappa$ maintained performance without compromising the known mAP of the PROB baseline.
Table S6: Ablation results for each hyperparameters of ETM and URM on OWOD split. The first row shows our default hyperparameter results. In subsequent rows, we vary a single hyperparameter, highlighted in green.
@@ C.5 Ablation Studies of each Module’s Hyperparameters
ETM URM Unknown Recall Known
$\delta$ $\sigma$ $k$ $\kappa$ Recall @1 mAP
1 1 10 0.5 30.56 11.69 58.89
1 1 10 0.45 31.42 11.98 58.38
1 1 10 0.55 30.11 11.47 59.12
1 1 5 0.5 28.51 10.58 59.31
1 1 15 0.5 32.01 11.89 58.80
1 0.5 10 0.5 31.55 10.69 59.02
1 2.0 10 0.5 31.52 11.59 58.95
0.9 1 10 0.5 30.70 11.20 58.89
1.1 1 10 0.5 31.68 11.93 58.79
Appendix D Additional Quantitative Results
D.1 Quantitative results on OWDETR split
Table S7 demonstrates that our method outperforms other models in both unknown object detection and feature learning on OWDETR split. This trend is consistent with the results on the Unknown-Unknown split shown in Table 2,3. The Unknown-Unknown split is a subset of the OWDETR split as described in B.1, which explains the similarity in performance trends.
Table S7: Results on OWDETR split. U-Recall denotes the unknown detection recall. Our method outperforms others in unknown detection and feature quality.
[…]
PROB 16.77 2.27 73.09
OSODD 16.77 2.22 73.09
UC-OWOD 10.62 1.27 22.84
RNCDL 17.09 1.52 72.34
PROB+URM 29.68 3.59 71.21
PROB+ETM 15.31 3.45 72.8
Ours 29.51 9.05 71.90
D.2 Quantitative Results of Efficiency Comparisons including Direct VFM usage
Table S8 reports runtime and memory costs, including a comparison with Ours+DINO, which pools DINOv2 features from proposals at inference. Our method introduces minimal overhead relative to PROB, requiring only a few additional MLP layers for instance feature extraction, and thus achieves comparable inference speed. In contrast, Ours+DINO Pool increases computation. The number of trainable parameters grows by only 0.5%, since SAM and DINO features are extracted once per image and reused during training.
We distill knowledge from VFMs rather than use them directly for two reasons: (1) VFMs are not designed for open-world detection, which requires identifying and incrementally learning unknown objects, and (2) direct use at inference is computationally costly.
Table S8: Efficiency comparisons including direct VFM usage. Our method enables effective feature learning with minimal computational overhead, compared to directly using VFM.
Metrics $\downarrow$ PROB Ours Ours+DINO Pool
Inference Time (msec) 47.1 47.2 80.9
Inference # Params 39.7M 39.9M 344M
Peak Inference Memory (MiB) 3030 3036 3036
Train Time (relative) x1.00 x1.74 x1.74
Peak Train Memory (MiB) 80660 80700 80700
Recall@1 4.42 11.69 15.53
D.3 Class Similarity Analysis